OpenStudy (anonymous):

What is the x-coordinate for the solution of the system of linear equations? 4x - 3y = -2 3x - y = -4

OpenStudy (anonymous):

y=3x+4 4x-3(3x+4)=-2 4x-9x-4=-2 -5x=2 x=-2/5

OpenStudy (anonymous):

Is -2/5 the answer??

OpenStudy (anonymous):

Yes